UsePoissonsheet to answer this question. Suppose that X, the number of customers arriving each ...

  1. UsePoissonsheet to answer this question.

Suppose that X, the number of customers arriving each hour at the only checkout counter in a local pharmacy, is approximately Poisson distributed with mean = 20 customers per hour. Find the probability that Xis more than 24. Round your answer to 4 decimal places. Put a box around your answer.  
Answer #1

\lambda = 20 customer per hour

P(x>24) = 1- P(<=24)

P(x>24) = 1 - Sum of Probabilities for x = 0 to 24)

Now we will use the poisson distribution table for lambda = 20.

Below is the the required probabilities -

P(x<=24) = 0+0+0+0+0+0.0001+0.0002+0.0005+0.0013+0.0029+0.0058+0.0106+0.0176+0.0271+0.0387+0.0516+0.0646+0.0760+0.0844+0.0888+0.0888+0.0846+0.0769+0.0669+0.0557

P(x<=24) = 0.8431

P(x>24) = 1 - P(x<=24)

P(x>24) = 1 - 0.8431

P(x>24) = 0.1569